# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 3.2.6. Calculations
The Floer theory on the U-shaped Lagrangian objects of $\mathcal{F}_{1}$ is the same as it was for the same U-shaped Lagrangians of $(Y,-v_{0})$ ([Can20, Lemma 5.14]),
@@
where $\partial$ is the differential, $C$ is a scalar function of $x$ defined by an open Gromov-Witten invariant from [KL15] and $\partial$ as a linear map is proportional to the linear map $\vartheta\otimes$- on $\operatorname{Ext}(\mathcal{L}^{i+1},\mathcal{L}^{j})$, the defining multi-theta function of $\Sigma_{2}$.
@@
Note that U-shaped Lagrangians go below the puncture, and wrap for large $r(t)$, so they do not interact with the puncture. The moduli space in Equation (3.12) admits a regular $J$ because there exist a dense set of such $J$ following the roadmap from [MS12] outlined in [Can20, §4.5, Lemma 4.47], and restricting the $J$ there to $Y^{0}$ here. Thus the moduli space counted in [Can20]
(3.31) $\mathcal{M}(p,q;\beta,J):=\left\{u:\mathbb{R}\times[0,1]\to Y\middle|\begin{array}[]{l}[u]=\beta,\ \overline{\partial}_{J}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{i},\ u(\mathbb{R},1)\subset L_{j},\\ u(-\infty)=q,\ u(+\infty)=p\end{array}\right\}/(u\sim u_{a})$
is isomorphic by a 1-1 bijection with the moduli space counted here (restricting $Y$ and $J$ to $Y_{0}$)
@@
Remark 3.21.
Because we are in the non-exact setting, changes in area affect calculations. In particular, homotoping $\gamma_{L_{j}}$ and $\delta_{K_{j}}$ (staying away from the critical values) to different choices then isotopes the Lagrangians above them and affects disc area calculations in the Fukaya category according to the formula [ACLL23a, Proposition 6.10]. However, in a symplectic Landau-Ginzburg model, the calculations split into fiber and base contributions because the Lagrangian isotopies (coming from parallel transport over homotopies) are exact [ACLL23b, Lemma 2.1], and different choices are therefore absorbed into $C$ by [ACLL23b, Corollary 2.4].
For morphisms between $\mathcal{F}_{1},\mathcal{F}_{2}$, suppose $\gamma_{L_{i}}\cap\delta_{K_{j}}$ intersect in two points as depicted in Figure 3 which are
$w_{1}:=\gamma_{L_{i}}(t_{1})=\delta_{K_{j}}(s_{1}),\quad w_{2}:=\gamma_{L_{i}}(t_{2})=\delta_{K_{j}}(s_{2}).$
There are bigons bounded by Lagrangians $L_{i}\cup K_{j}$ covering the bigon in the base between these two paths and intersection points. Hamiltonian-perturb that bigon to a circle. That is, let
$\zeta:S^{1}\to\mathbb{C}^{*}$
denote the projection of a portion of Hamiltonian-equivalent Lagrangians $\phi_{H}(L_{i}\cup K_{j})$ which project to a circle centered at $T^{\epsilon}$ and let $\zeta(1)$ be the image of $w_{1}$ under the Hamiltonian diffeomorphism composed with $W_{0}$. Then define (where subscripts $Y_{w}:=Y|_{W_{0}^{-1}(w)}$ denote the fiber over that point)
$\Phi^{s}:Y^{0}_{w_{1}}\to Y^{0}_{\zeta(1)}\to Y^{0}_{\zeta(e^{2\pi is})}\to Y^{0}_{\phi_{H}^{-1}(\zeta(e^{2\pi is}))}$
to be the composition of the Hamiltonian isotopy, parallel transport counterclockwise around the circle an angle of $s$ from the first intersection point at $w_{1}$, and then reversing the Hamiltonian isotopy back to the original Lagrangians. Lastly, let
$\ell_{i}^{\prime}:=\Phi_{\gamma_{L_{i}}(t_{1})}(\ell_{i}),\qquad\ell_{j}^{\prime\prime}:=\Phi_{\delta_{K_{j}}(s_{1})}(\ell_{j}),\qquad\Phi:=\Phi^{1}$
so $\Phi=\Phi^{1}$ denotes the monodromy of $W_{0}$ around the singular fiber. With this set-up, we can now examine intersections between $L_{i}$ and $K_{j}$.
@@
$CF(L_{i},K_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$.
@@
(3.33) $CF_{Y^{0}}(L_{i},K_{j})$ $=HF_{Y^{0}_{w_{1}}}(\ell^{\prime}_{i},\ell^{\prime\prime}_{j})[1]\oplus HF_{Y^{0}_{w_{2}}}(\Phi^{1-s}(\ell^{\prime}_{i}),\Phi^{-s}(\ell^{\prime\prime}_{j}))$
$=HF_{T^{4}}(\ell_{i},\ell_{j})[1]\oplus HF_{T^{4}}(\Phi^{1}(\ell_{i}),\ell_{j})$
$=HF_{T^{4}}(\ell_{i},\ell_{j})[1]\oplus HF_{T^{4}}(\ell_{i+1},\ell_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$
where we use $HF$ on the righthand-side because in a fiber $\partial=0$ as there are no bigons between linear Lagrangians in a 4-torus. We start with $\ell^{\prime}_{i},\ell^{\prime\prime}_{j}$ in the fiber over the left intersection point, and parallel transport around the loop around the origin with respect to that starting point. The second equality follows because $\Phi^{s}$ is a diffeomorphism for any $s$ so preserves intersection points, hence we can apply $\Phi^{s}$ to Lagrangians in the second factor of the $\oplus$ and then note that $\Phi^{1}(\ell_{i})$ is Hamiltonian isotopic to $\ell_{i+1}$ [Can20, Lemma 4.24] so the statement follows from the Hamiltonian invariance of $HF$ [Aur14, Theorem 1.5]. The gradings will be justified in Subsection 4.1.
@@
[…]
$L_{i}$
$\phi(L_{i})$
@@
(a) The base as a cylinder. The Lagrangian intersection points of $CF(L_{i},K_{j})$ are denoted in the two fibers over the base intersections, upon positively wrapping the ends of $L_{i}$. Here from the left intersection point to the right, we go counterclockwise $2\pi(1-s)$ around $W_{0}(\phi(L_{i}))$ and clockwise $2\pi s$ around from $W_{0}(K_{j})$. So up to diffeomorphism the differential counts bigons between two fibers, $\ell_{i}\cap\ell_{j}$ and $\Phi(\ell_{i})\cap\ell_{j}$ where $\Phi$ denotes monodromy around the singularity, and $\Phi(\ell_{i})\cong\ell_{i+1}$.
@@
[…]
$L_{i}$
$\phi(L_{i})$
@@
Figure 3. Projection of $K_{i}$, $L_{j}$, and the positively wrapped $\phi(L_{j})$ to the base of $W_{0}:Y^{0}\to\mathbb{C}$. There is a puncture at 0, and the singular fiber is at $T^{\epsilon}$, denoted by $\times$.
Now we compute the differential and pass to cohomology.
Lemma 3.23 ($\operatorname{Hom}(\mathcal{F}_{1},\mathcal{F}_{2})$).
Recall $CF(L_{i},K_{j})\cong CF(\ell_{i},\ell_{j})[1]\oplus CF(\ell_{i+1},\ell_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$. The differential can be computed as in [Can20, Lemma 5.14] so Floer groups are again defined from
$CF(\ell_{i+1},\ell_{j})$ $\xrightarrow{\partial}CF(\ell_{i},\ell_{j})[1]\to HF(L_{i},K_{j})\to 0$ to be
$HF(L_{i},K_{j})=\mathbb{C}^{(i-j)^{2}}/C\partial(\mathbb{C}^{(i+1-j)^{2}}).$
@@
Since the moduli spaces are isomorphic, we obtain the same count as in Lemma 3.20 which was for two U-shaped curves. An illustration of how the two moduli spaces can be identified is given in Figure 4. ∎
@@
[…]
$\phi(L_{i})$
@@
Figure 4. The differential calculation between one U-shape and one arc on $Y^{0}$ is equivalent to the differential calculation between two U-shapes on $Y$ by including $Y^{0}\hookrightarrow Y$, applying a Lagrangian isotopy $\phi^{\prime}$ on the arc $K_{j}$ by homotoping the curve in the base as shown in solid line, and then completing it to a U-shape as shown with the additional dotted line.
Lemma 3.24 ($\operatorname{Hom}(\mathcal{F}_{2},\mathcal{F}_{1})$).
@@
Morphisms are defined by positive isotopies of the input Lagrangian that wrap arbitrarily close to $\pi$ around $\infty$ in the $\mathbb{C}^{*}$ model of the base by Equation (3.29). Thus the end of the projected input Lagrangian will wrap past the ends of the projected output Lagrangian. In other words, we displace $K_{j}$ from $L_{i}$ in the base so that it is a ray that lies entirely above $L_{i}$ in the base. Since it is already entirely above $L_{i}$ to start, there are no intersection points in the base so there can be no intersection points in the total space. ∎
This exhibits the semi-orthogonality between $\mathcal{F}_{1},\mathcal{F}_{2}$ because there are morphisms in one direction and none in the other.
Lemma 3.25 (Morphisms on $\mathcal{F}_{2}$).
$HF(K_{i},K_{j})\cong\mathbb{C}^{(j-i)^{2}}[x].$
@@
Intersection points in the total space cover intersections in the base. By Equation (3.3), we only fully wrap at one end of the cylinder. Counting intersections and triangles in the base cylinder (a standard calculation [Abo10], [Aur14, §4.2]) we find that $HF(W_{0}(K_{i}),W_{0}(K_{j}))\cong\mathbb{C}[x]$ as algebras. In the torus fiber $CF(\ell_{i},\ell_{j})=HF(\ell_{i},\ell_{j})=\mathbb{C}^{(j-i)^{2}}$ ([Fuk02], [ACLL, §4.2]). Thus over each intersection point corresponding to $x^{k}$ in the base we have $\mathbb{C}^{(j-i)^{2}}$ intersection points in the fiber and
$HF(K_{i},K_{j})=CF(\ell_{i},\ell_{j})\otimes\mathbb{C}[x]\cong\mathbb{C}^{(j-i)^{2}}[x].$
4. Product structure
We compose from inside to out, meaning from right to left. Due to orthogonality, composition on the B-side is nonzero in two cases not covered above; either the first morphism is on the $H$ factor and then we compose with a morphism from $H$ to the $V\times\mathbb{C}$ factor, or the first morphism is from the $H$ factor to the $V\times\mathbb{C}$ factor and we compose with one on $V\times\mathbb{C}$. In other words,
$\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))\otimes\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)))$
$\hskip 72.26999pt\to\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))$

### v2

@@ 3.4. Calculations of morphisms
The Floer theory on the U-shaped Lagrangian objects of $\mathcal{A}_{L}$ is isomorphic to the Floer theory of the U-shaped Lagrangians of $(Y,-v_{0})$ in [Ca20, Lemma 5.14]. That is,
@@
where $\partial$ is the differential between Floer complexes $CF(\ell_{i+1},\ell_{j})[-1]\cong\mathbb{C}^{(i+1-j)^{2}}\to CF(\ell_{i},\ell_{j})\cong\mathbb{C}^{(i-j)^{2}}$ and $C$ is a scalar function of $(x_{1},x_{2},y)$ defined by an open Gromov-Witten invariant from [KL15].
In particular [Ca20] shows that $\partial$, as a linear map, is proportional to the linear map $\vartheta\otimes$- on $\operatorname{Ext}(\mathcal{L}^{i+1},\mathcal{L}^{j})\to\operatorname{Ext}(\mathcal{L}^{i},\mathcal{L}^{j})$, for $\vartheta$ the defining theta function of $H=\Sigma_{2}$, thus proving homological mirror symmetry of morphisms that $\operatorname{Ext}^{*}(\mathcal{L}^{i}|_{H},\mathcal{L}^{j}|_{H})=HF^{*}(L_{i},L_{j})$.
@@
[…]
Figure 5. The differential on $CF(L_{i},L_{j})$ counts curves covering this bigon depicted in the base, from point $b_{1}$ of degree 1-1 to point $b_{0}$ of degree 0. : Add $\tfrac{1}{4}\cdot 2\pi$ as the oriented tangent vector rotates counterclockwise $\frac{\pi}{2}$ from $b_{0}$ to $b_{1}$
The differential on $CF(L_{i},L_{j})$ counts curves covering the bigon depicted in the base in Figure 5. Note that U-shaped Lagrangians stay outside the $R_{2}$ disc by Definition 3.3, so they do not interact with the puncture at 0. The moduli space in Equation (3.19) admits a regular $J$ as follows. There exists a dense set of regular $J$ on $Y$ following the roadmap from [jholbk] outlined in [Ca20, §4.5, Lemma 4.47] for the moduli space
(3.58) $\mathcal{M}(p,q;\beta,J):=\left\{u:\mathbb{R}\times[0,1]\to Y\middle|\begin{array}[]{l}[u]=\beta,\ \overline{\partial}_{J}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{i},\ u(\mathbb{R},1)\subset L_{j},\\ u(-\infty)=q,\ u(+\infty)=p\end{array}\right\}/(u\sim u_{a}).$
Then we restrict $J$ to $Y^{0}$ and note that the moduli space of Equation (3.58) is isomorphic by a 1-1 bijection with the moduli space counted in the differential here
@@
Then we carry out the same calculation of the differential as in [Ca20, Lemma 5.14].
Because we are in the non-exact setting, changes in area affect calculations. In particular, homotoping $\gamma_{L_{j}}$ and $\delta_{K_{j}}$ (staying away from the critical values) to different choices then isotopes the Lagrangians above them by parallel transport and affects disc area calculations in the Fukaya category according to the formula [ACLL_Bfield, Proposition 6.10]. However, in a symplectic Landau-Ginzburg model, the calculations split into fiber and base contributions because the Lagrangian isotopies are exact [ACLL_LG, Lemma 2.1]. Therefore, different choices can be absorbed into $C$ by [ACLL_LG, Corollary 2.4]. ∎
For morphisms from $\mathcal{A}_{L}$ to $\mathcal{A}_{K}$, recall that we perturb the ends of the input projected Lagrangian to lie above those of the output, Equation (3.42). After wrapping, set the notation that ${L_{i}}\cap{K_{j}}$ intersect in two points $w_{0},w_{1}$ over points $b_{0},b_{1}\in\gamma_{L_{i}}\cap\delta_{K_{j}}$ in the base as depicted in Figure 6 which are
$b_{0}=W_{0}(w_{0}):=\gamma_{L_{i}}(t_{0})=\delta_{K_{j}}(s_{0}),\quad b_{1}=W_{0}(w_{1}):=\gamma_{L_{i}}(t_{1})=\delta_{K_{j}}(s_{1}).$
The differential counts bigons bounded by Lagrangians $L_{i}\cup K_{j}$ that cover the bigon in the base between these two paths and intersection points. Hamiltonian-perturb the Lagrangians via Hamiltonian $\tilde{H}$, so that the boundary of the bigon in the base, $W_{0}(\phi_{\tilde{H}}(L_{i}\cup K_{j}))$, becomes a circle centered at $T^{\epsilon}$ described by
$\zeta:S^{1}\to\mathbb{C}^{*}.$
Parametrize $\zeta$ so that it starts at $\zeta(1)=W_{0}(\phi_{\tilde{H}}(w_{0}))$. Setting the notation $Y^{0}_{b}:={W_{0}^{-1}(b)}$ to denote the fiber over $b$, define
$\Phi^{s}:Y^{0}_{b_{0}}\to Y^{0}_{\zeta(1)}\to Y^{0}_{\zeta(e^{2\pi is})}\to Y^{0}_{b_{0}^{\prime}}$
to be the composition of the Hamiltonian isotopy on $Y^{0}_{b_{0}}$, parallel transport counterclockwise around the circle an angle of $2\pi s$, and then the inverse of the Hamiltonian isotopy, bringing us to a fiber $Y_{b_{0}^{\prime}}$. Lastly, let
$\ell_{i}^{\prime}:=\Phi_{\gamma_{L_{i}}(t_{0})}(\ell_{i})=L_{i}|_{b_{0}},\qquad\ell_{j}^{\prime\prime}:=\Phi_{\delta_{K_{j}}(s_{0})}(\ell_{j})=K_{j}|_{b_{0}},\qquad\Phi:=\Phi^{1}$
so $\Phi=\Phi^{1}$ denotes the monodromy of $W_{0}$ around the singular fiber. With this set-up, we can now examine Floer theory between $L_{i}$ and $K_{j}$. We take the convention on cochain complexes that $C^{n}[p]=C^{n+p}$ and consider Floer theory with $K_{j}[2]$, which has phase 0 by Equation (3.50). With the choice of grading on $L_{i},K_{j}[2]$ in Definition 3.23, the degrees of the base intersection points matches that in Figure 4, so appearing in degrees 0 and 1.
@@
$CF^{*}(L_{i},K_{j}[2])\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$.
@@
(3.60) $CF_{Y^{0}}^{*}(L_{i},K_{j}[2])$ $=HF^{*}_{Y^{0}_{b_{0}}}(\Phi_{\gamma_{L_{i}}(t_{0})}(\ell_{i}),\Phi_{\delta_{K_{j}}(s_{0})}(\ell_{j}))[1]\oplus HF^{*}_{Y^{0}_{b_{1}}}(\Phi^{1-s}(\ell^{\prime}_{i}),\Phi^{-s}(\ell^{\prime\prime}_{j}))$
$=HF^{*}_{Y^{0}_{b_{S}}}(\Phi_{\delta_{K_{j}}(s_{0})}^{-1}\Phi_{\gamma_{L_{i}}(t_{0})}(\ell_{i}),\ell_{j})[1]\oplus HF^{*}_{Y^{0}_{b_{1}}}(\Phi^{1}(\ell_{i}^{\prime}),\ell_{j}^{\prime})$
$=HF^{*}_{T^{4}}(\ell_{i},\ell_{j})[1]\oplus HF^{*}_{T^{4}}(\ell_{i+1},\ell_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$
where we use $HF$ on the righthand-side because in a fiber $\partial=0$ as there are no bigons between linear Lagrangians in a 4-torus. We start with $\ell^{\prime}_{i},\ell^{\prime\prime}_{j}$ in the fiber over the left intersection point $b_{0}$, and parallel transport around the singular fiber. The second equality follows because parallel transport $\Phi_{\gamma(0)\to\gamma(1)}$ is a diffeomorphism so preserves intersection points, hence we can apply $\Phi_{\delta_{K_{j}}(s_{0})}^{-1}$ and $\Phi^{s}$ to Lagrangians in the first and second factor of the direct sum, respectively. Then note that $\Phi_{\delta_{K_{j}}(s_{0})}^{-1}\Phi_{\gamma_{L_{i}}(t_{0})}$ is parallel transport around a loop not containing the critical value so is Hamiltonian by [symp_intro, Theorem 6.21(ii)], and $\Phi^{1}(\ell_{i})$ is Hamiltonian isotopic to $\ell_{i+1}$ by [Ca20, Lemma 4.24]. Thus the statement follows from the Hamiltonian invariance of $HF$ [fuk_intro, Theorem 1.5].
@@
$\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}W_{0}(K_{j}[2])$}
$W_{0}(L_{i})$
$W_{0}(\phi(L_{i}))$
@@
$b_{0}$
$b_{1}$
(a) The base as a cylinder. The Lagrangian intersection points of $CF^{*}(L_{i},K_{j}[2])$ lie in two fibers over the base intersections, upon positively wrapping the ends of $L_{i}$ by $\phi$. Here from the left intersection point to the right, we move counterclockwise $2\pi(1-s)$ around $W_{0}(\phi(L_{i}))$ and clockwise/to the right $2\pi s$ along $W_{0}(K_{j}[2])$. So up to diffeomorphism the differential counts bigons between two fibers, $\ell_{i}\cap\ell_{j}$ and $\Phi(\ell_{i})\cap\ell_{j}$ where $\Phi$ denotes monodromy around the singularity, and $\Phi(\ell_{i})\cong\ell_{i+1}$ in the Fukaya category.
@@
$\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}W_{0}(K_{j}[2])$}
$W_{0}(L_{i})$
$W_{0}(\phi(L_{i}))$
@@
Figure 6. Projection of $K_{i}[2]$, $L_{j}$, and the positively wrapped $\phi(L_{j})$ to the base of $W_{0}:Y^{0}\to\mathbb{C}$. There is a puncture at 0, and the singular fiber is at $T^{\epsilon}$, denoted by $\times$.
Now we compute the differential and pass to cohomology. Note that $HF(L_{i},K_{j}[2])$ is supported in degree 1 but the mirror morphism group is supported in degree 0. We take a shift of $K_{j}[2]$ down by 1, accordingly, and then $K_{j}[1]$ has phase 1/2-1/2 in the base, analogous to the U-shape ends.
Lemma 3.27 ($\operatorname{Hom}(\mathcal{A}_{L},\mathcal{A}_{K})$).
The cohomology of $CF(\ell_{i+1},\ell_{j})[-1]$ $\xrightarrow{\partial}CF(\ell_{i},\ell_{j})\to HF(L_{i},K_{j}[1])\to 0$ is
(3.61) $HF^{*}(L_{i},K_{j}[1])=HF^{*}(L_{i},L_{j})=\mathbb{C}^{(i-j)^{2}}/C\partial(\mathbb{C}^{(i+1-j)^{2}}).$
@@
The moduli spaces counted in the differential are isomorphic to those counted for two U-shapes in Lemma 3.25, as illustrated in the identification in Figure 7. Note that the grading on $K_{j}[1]$, under this identification, matches with the grading on the U-shaped Lagrangian $L_{j}$.∎
@@
0
$\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}W_{0}(K_{j}[1])$}
$W_{0}(\phi(L_{i}))$
@@
Figure 7. The differential on $CF(L_{i},K_{j}[1])$ is equivalent to the differential on $CF(L_{i},L_{j})$. This can be seen by applying a Lagrangian isotopy $\phi^{\prime}$ on the arc $K_{j}[1]$ homotoping the curve in the base as shown in the solid line, and then completing it to a U-shape as shown with the additional dotted line. The resulting moduli spaces of bigons in the differentials can then be identified. Furthermore, the gradings on $K_{j}[1]$ and $L_{j}$ agree under this identification near the intersection points.
Lemma 3.28 ($\operatorname{Hom}(\mathcal{A}_{K},\mathcal{A}_{L})$).
@@
Recall from Equation (3.42) that we first wrap the ends of the projected output Lagrangian negatively arbitrarily close to $\pi$, or equivalently we can wrap that of the input $W_{0}(K_{j})$ positively arbitrarily close to $\pi$. Then chain-level morphisms are given by the cochain complex defined in Equation (3.5) generated by intersection points between the resulting wrapped input and original output. Wrapping here does not introduce more intersection points because the end of the projected input Lagrangian $W_{0}(K_{j})$ already lies above (in other words, counterclockwise from) the ends of the projected output Lagrangian $W_{0}(L_{i})$, see Figure 6. No intersection points in the base means no intersection points in the total space and therefore the morphism group $HF(K_{j},L_{i})=0$. ∎
This exhibits the semi-orthogonality between $\mathcal{A}_{L},\mathcal{A}_{K}$ because there are morphisms in one direction and none in the other.
Lemma 3.29 (Morphisms on $\mathcal{A}_{K}$).
(3.62) $HF^{d}(K_{i},K_{j})\cong\begin{cases}\mathbb{C}^{(j-i)^{2}}[x]&j>i,d=0\text{ or }j<i,d=2\\ \mathbb{C}^{d\choose 2}[x]&i=j\\ 0&\text{ else }\end{cases}.$
@@
Intersection points in the total space cover intersections in the base of $W_{0}$. Morphisms and their composition in the base cylinder are given by
(3.63) $HF^{*}_{\mathbb{C}^{*}}(W_{0}(K_{i}),W_{0}(K_{j}))\cong\mathbb{C}[x]$
as algebras, supported in degree 0, by counting intersections and triangles in the universal cover, using a quadratic Hamiltonian as explained in Remark 3.7. Note that the differential is 0 in the base because there is no bigon between two lines of fixed slope. The details of the calculation are given in [fuk_intro, §4.2], using a rescaling trick. Note that the coordinates $r,\theta$ in [fuk_intro, §4.2] are $\mu,\theta$ here.
In a torus fiber over a base intersection point, $CF_{T^{4}}^{*}(\ell_{i},\ell_{j})=HF^{*}_{T^{4}}(\ell_{i},\ell_{j})$ is given by Equation (3.53). Thus over each base intersection point corresponding to $x^{k}$ under the isomorphism in Equation (3.63), we have
(3.64) $HF^{d}(K_{i},K_{j})=CF^{d}(\ell_{i},\ell_{j})\otimes\mathbb{C}[x]\cong\begin{cases}\mathbb{C}^{(j-i)^{2}}[x]&j>i,d=0\text{ or }j<i,d=2\\ \mathbb{C}^{d\choose 2}[x]&i=j\\ 0&\text{ else }\end{cases}.$
Note that by [Ab10], $W_{0}(K_{j})$ split-generates the fully-wrapped Fukaya category of $T^{*}S^{1}$.
3.5. Product structure
We compose morphisms from inside to out, meaning from right to left. Due to orthogonality on the A- and B-sides, composition between factors is nonzero in two cases.
B-side: Composition in the bounded derived category of coherent sheaves reduces to multiplication of theta functions, restricted to the hypersurface.
First, we have composition within the $H$ factor and also within the $V\times\mathbb{C}$ factor. Between factors, we can express the two possible cases roughly as
$\text{Hom}_{D^{b}(X)}(D^{b}(H),D^{b}(V\times\mathbb{C}))\otimes\text{Hom}_{D^{b}(X)}(D^{b}(H),D^{b}(H))\to\text{Hom}_{D^{b}(X)}(D^{b}(H),D^{b}(V\times\mathbb{C}))$